OpenStudy (anonymous):

A hemispherical vessel is partly filled with water 12 cm deep at the lower part and oil 6 cm deep above the water. If the volume of the water is equal to the volume of the oil. What is the Volume of the oil

OpenStudy (alekos):

whats the radius of the hemisphere?

OpenStudy (anonymous):

It is not given

OpenStudy (alekos):

so we have to assume that the vessel is full?

OpenStudy (anonymous):

I think so

OpenStudy (alekos):

so the volume is 1/3(pi*h^2)(3r-h) where r=18 and h=12

OpenStudy (anonymous):

The answer to the question should be 8143 cu.cm

OpenStudy (anonymous):

And the radius should be 22. But I don't know how to get 22 as the radius

OpenStudy (alekos):

you need the radius to work it out in the first place

OpenStudy (hoblos):

 the volume of a spherical cap is \[V= \frac{ \pi \times h }{ 6 }(3a ^{2}+h ^{2})\] volume of water = pi*12(3a^2 + 12^2)/6 volume of oil = volume of both oil and water - volume of water = pi*(12+6)(3b^2 + 18^2)/6 - pi*12(3a^2 + 12^2)/6 if volume of oil = volume of water then pi*(12+6)(3b^2 + 18^2)/6 - pi*12(3a^2 + 12^2)/6 = pi*12(3a^2 + 12^2)/6 pi*(12+6)(3b^2 + 18^2)/6 = 0 so now you can solve for b once you find b you can get the total volume of both oil and water which is pi*(12+6)(3b^2 + 18^2)/6 and since the two volumes are equal the volume of oil would be half the volume of both